OpenStudy (anonymous):

convert -140 degrees to radians

OpenStudy (anonymous):

use degreesx(pi/180)

OpenStudy (anonymous):

140(pi/180)=-7pi/9

OpenStudy (anonymous):

-140(pi/180)=-7pi/9

OpenStudy (amistre64):

degrees need pi; *pi/180

OpenStudy (anonymous):

- 140 degrees is the same as 360 + -140.. So that would be 220 degrees. And to convert to radians you do \[220 * \frac{\pi}{180}\] which is \[\frac{220\pi}{180}\] or \[\frac{11\pi}{9}\]